OpenStudy (anonymous):

Crystal can wax a floor in 16 minutes. One day her friend perry helped her and it only took 5.76 minutes. How long would it take perry to do it alone?

OpenStudy (anonymous):

in 1 min prerry can do=1/5.76-1/16=0.111111111 so perry alone can do =1/0.111111111=9 min

OpenStudy (anonymous):

yes i get it!